# Towards Efficient Real-Time Video Motion Transfer via Generative Time Series Modeling

### v1

@@ Abstract
We propose a deep learning framework designed to significantly optimize bandwidth for motion-transfer-enabled video applications, including video conferencing, virtual reality interactions, health monitoring systems, and vision-based real-time anomaly detection. To capture complex motion effectively, we utilize the First Order Motion Model (FOMM), which encodes dynamic objects by detecting keypoints and their associated local affine transformations. These keypoints are identified using a self-supervised keypoint detector and arranged into a time series corresponding to the successive frames. Forecasting is performed on these keypoints by integrating two advanced generative time series models into the motion transfer pipeline, namely the Variational Recurrent Neural Network (VRNN) and the Gated Recurrent Unit with Normalizing Flow (GRU-NF). The predicted keypoints are subsequently synthesized into realistic video frames using an optical flow estimator paired with a generator network, thereby facilitating accurate video forecasting and enabling efficient, low-frame-rate video transmission. We validate our results across three datasets for video animation and reconstruction using the following metrics: Mean Absolute Error (MAE), Joint Embedding Predictive Architecture Embedding Distance (JEDi), Structural Similarity Index (SSIM), and Average Pair-wise Displacement (APD). Our results confirm that by utilizing the superior reconstruction property of the Variational Autoencoder, the VRNN integrated FOMM excels in applications involving multi-step ahead forecasts such as video conferencing. On the other hand, by leveraging the Normalizing Flow architecture for exact likelihood estimation, and enabling efficient latent space sampling, the GRU-NF based FOMM exhibits superior capabilities for producing diverse future samples while maintaining high visual quality for tasks like real-time video-based anomaly detection.
@@ 1 Introduction
With the growing demand for ubiquitous services, telecommuting, immersive visual interactions, and realistic simulations, applications such as video conferencing, augmented reality (AR), virtual reality (VR) gaming, and remote medical monitoring are becoming increasingly widespread [4, 62, 68]. These applications demand not only high-quality visual outputs and seamless user interaction but also low-latency processing and efficient bandwidth usage—challenges that conventional video streaming and compression techniques struggle to meet effectively. Video motion transfer [62] presents an effective and promising approach for realizing such applications by transferring motion from a driving video to a source frame or video, thereby enhancing user experiences and enabling more creative and engaging interactions. A key challenge in this space lies in modeling and predicting fine-grained motion or object dynamics in a way that preserves both temporal coherence and visual fidelity, while also being bandwidth-efficient. Pixel-based methods if used for prediction can suffer from high computational cost, error accumulation, and limited generalization in dynamic and diverse environments [37]. Recent advances in representation learning and generative time series modeling have started to address these limitations by exploring latent spaces, structured motion representations, and probabilistic forecasting [37]. Among various representation learning based techniques for video synthesis and prediction, keypoint-based methods stand out by extracting keypoints from source and driving videos for enabling precise motion transfer [50, 51] while achieving superior video quality and compression beyond traditional codec-based methods. This not only reduces computational and bandwidth overhead, but also enables fine-grained control over motion dynamics, making keypoint based methods ideal for real-time applications as compared to competing approaches which also claim to perform motion synthesis with high fidelity [39]. A critical advantage of using keypoint based motion transfer architectures is their ability to provide a natural structure for incorporating temporal prediction using generative time series models [7, 42] enabling generalization across a variety of subjects and contexts.
Real-time motion transfer based applications using keypoint based representations can involve multi-step ahead video forecasts for applications such as telecommuting as well as diverse future video prediction for applications like vision-based anomaly detection. Some examples of real-time keypoint-based motion transfer applications are described as follows:
@@
  • Video Conferencing: In streaming applications such as video conferencing, keypoint-based motion transfer enables efficient future frame prediction by transmitting keypoints of initial frames from a source mobile or client, allowing the pipeline to predict and generate future frames at the receiving end [62]. Such an architecture not only reduces the dependency on ubiquitous network connectivity, but also optimizes computational efficiency and bandwidth utilization.
@@
  • Telehealth: For certain critical healthcare conditions it is essential for healthcare providers along with supporting personnel such as interns to monitor patients remotely. However, in such cases patients may be hesitant to reveal their identity in virtual settings [68] and this could also be mandated by regulatory restrictions. Keypoint-based motion transfer enables future frame forecasting with anonymization that ensures both data privacy and accurate transmission of critical body movements of the patient for effective remote monitoring.
@@
  • Manufacturing Industry: A key technical challenge when deploying Artificial Intelligence (AI)-driven inspection systems in manufacturing is the scarcity of comprehensive real-world datasets that include diverse defect types. Such limitations hinder the effectiveness of machine learning models, potentially resulting in missed detections or false negatives if predictions do not correspond precisely to known defects. In vision-based anomaly detection, by forecasting future video frames of units on the production line, the system can anticipate and visualize the onset of potential defects in upcoming frames. Generating multiple plausible future sequences instead of a single prediction increases the chances of capturing visual patterns that may correspond to defects. This approach facilitates the generation of early and reliable defect warnings during real-time inspections. Moreover, the synthetic defect datasets generated through this approach expand the training data available for downstream classification or defect detetion algorithms, enhancing their exposure to a wider variety of defect scenarios and improving overall robustness and detection accuracy.
@@
In this work we go beyond pre-existing keypoint based motion transfer frameworks such as the First Order Motion Model (FOMM) [51] which provide spatial compression and focus on integrating probabilistic sequential generative models for keypoint forecasting in the motion transfer pipeline for realizing both spatial and temporal compression. In previous work [3], this has been demonstrated for applications involving video prediction using a Variational Recurrent Neural Network (VRNN) integrated within the FOMM pipeline. In our current work we substantially expand on these preliminary findings and also investigate other generative models including Normalizing Flow and Gated Recurrent Unit-Normalizing Flow (GRU-NF) for keypoint forecasting within the motion transfer pipeline for both video prediction as well as diverse sample generation tasks. In the latter case we perform a systematic study of video diversity versus quality using a variety of metrics for measuring differences in generated realizations as well as their perceptual quality. Our enhanced architecture in this case can provide upto 20x bandwidth savings which is over the 10x savings that can be realized from existing keypoint based motion transfer pipelines for video reconstruction and animation [50, 51, 62, 72]. To the best of our knowledge our work is the first systematic exploration of the potential of Normalizing Flow (NF) based architectures for predictive inference within a keypoint based motion transfer framework.
@@
The rest of the paper is organized as follows. Section 2 discusses related works on video prediction, motion transfer and diverse sample generation. Section 3 describes our proposed real-time motion transfer architecture and the different forecasting models used in our work. The evaluation metrics used in this work to assess the performance of the models are explained in Section 4. Section 5 analyzes the results of our simulations based on the evaluation metrics for estimating the accuracy of generated videos and provides a comparative analysis of the keypoint predictor models used. Finally, in Section 6 we present our conclusions and our plans for future work.
@@ 2 Related Works
Single Sample Video Prediction: The prediction and synthesis of videos that accurately represent dynamic object movements is crucial for numerous practical applications. For this purpose various video prediction techniques including direct pixel synthesis, narrowing the prediction space to high-level feature spaces [37] have been explored by researchers. In order to accurately predict and synthesize videos by modeling object dynamics at the pixel level, numerous feature learning strategies have been explored, including adversarial training [29] and gradient difference loss functions [32]. Various recurrent and convolutional networks such as Convolutional Long-Short Term Memory (ConvLSTM) [47], E3d-LSTM [63], and MSPred [57] are used for future frame prediction in raw pixel space. The common challenge faced by recurrent networks is vanishing gradient that hinders them to forecast in long term [15]. To address this, transformer-based models incorporating self-attention mechanisms have been explored for long-term video prediction [48, 69, 28]. While transformers have demonstrated strong performance on large text and image datasets, they are not yet well-optimized for fine-grained spatial-temporal modeling [35]. To mitigate these challenges, a simpler video prediction model entirely based on CNNs, which outperformed both recurrent neural networks (RNNs) and Vision Transformers (ViTs) has been proposed in [15]. The use of CNNs in video frame prediction has been explored by creating different types of models, i.e., GDL [32], PredCNN [65], and SDC-Net [43]. However, CNN-based frame synthesis is not suitable for spatially shifting pixels in input frames and is also inefficient for forecasting in large data sets [35]. Generative Adverserial Network (GAN) models have also been used to predict a sequence of future frames based on a sequence of initial video frames using pixel space as input [1, 21]. Additionally, flow-based architectures have been explored as an alternative that enables exact maximum likelihood learning utilizing invertible transformations. To perform forecasting, latent embeddings of video frames instead of raw pixels are used and future frames are reconstructed through an invertible mapping in [40]. The key finding here is that NF based models outperform GANs that are based on adversarial training.
@@
Inspite of their proliferation, pixel-based techniques face considerable challenges, especially for tasks requiring precise future predictions involving small or slow-moving objects. In addition, pixel-level prediction methods commonly accumulate significant errors due to inherent variability between consecutive frames. To address these challenges associated with high-dimensional pixel spaces, researchers have increasingly turned towards representation learning methods, such as semantic segmentation, human pose estimation, and keypoint-based representations. Before the widespread adoption of deep neural networks, traditional approaches employed Principal Component Analysis (PCA) or pose parameters for reconstructing video sequences with limited data, such as face boundaries [27, 24]. Leveraging the combined strengths of VAEs and GANs, Walker et al. [60] proposed an architecture that uses VAE to model possible future human poses in pixel space and then predicting future frames using a GAN with future poses used as conditional information. However, GANs often produce blurry and temporally inconsistent outputs [35]. Similarly, VAEs struggle to model complex video dynamics as a result of collapsing of posterior distribution to the prior [73]. Overall, pose-guided prediction methods, despite showing promise, have largely been restricted to videos containing human subjects [53, 58, 59, 60]. Semantic segmentation is another representation learning technique which instead of using raw pixel data extracts semantic elements by segmenting current frames and predict future optical flows through temporal models like Long Short-Term Memory (LSTM) networks [41, 54, 19]. Keypoint-based approaches for representation learning, on the other hand, explicitly represent object-level dynamics, and can provide superior outcomes for trajectory prediction and action recognition. Recent research has demonstrated the efficacy of keypoints combined with reference frames for reconstructing future video frames, thereby substantially reducing accumulated errors inherent in pixel-space forecasting [34]. In the context of using latent space embeddings, some recent works tackle the prediction challenge by using advanced modeling techniques instead of relying on basic linear motion predictions derived from past frames [35]. In [13], a CNN-based vector quantization approach is introduced that discretizes images into a compact latent representation, enabling transformer-based autoregressive modeling for high-resolution image synthesis. Compressing high resolution videos into latent variables by using a vector quantized variational autoencoder (VQ-VAE) allowing for high resolution predictions has been proposed in [61]. However, both the CNN and VQ-VAE based models often require extensive computational resources and substantial training data. Taking these limitations into account, our work employs keypoint-based prediction methodologies, integrating these representations effectively into video reconstruction and animation in a real-time motion transfer pipeline across multiple diverse datasets.
@@
Diverse Sample Video Prediction: Incorporating uncertainty is another category of video prediction methods that allows generating diverse outcomes from a single input [37]. Video frames form multi-modal distributions in a high dimensional space and given this level of complexity, a range of approaches have incorporated latent variables into deterministic models e.g., LSTM and CNN to introduce stochasticity and variation in predictions. Gaussian Processes combined with LSTM have been used to enable the generation of diverse samples by encoding prior knowledge of future states from past observations, allowing for a more structured distribution over possible outcomes [49]. Latent variables have also been introduced into convolutional autoencoders to model video unpredictability through generative functions that capture diverse frame reconstructions [16]. Similarly, convolutional recurrent networks have been trained to represent future stochasticity and generate varied sequences by sampling from learned latent distributions [20]. Stochastic neural networks extend deterministic models by integrating Gaussian latent variables with convolutional encoders and spatial transformers, enabling the synthesis of multimodal futures with enhanced diversity [14]. The SV2P framework advances this direction by extending the CDNA model to include latent variables, allowing for different possible futures per input sample [2]. SVG further improves uncertainty modeling by introducing a time-varying prior within an LSTM autoencoder, which better captures temporal dependencies in video sequences [9].
Another common approach for generating diverse predictions is by sampling from the latent space of generative models such as VAE, GAN, and NF that can learn rich probabilistic representations of data [37]. To explore diverse video frame outcomes, GANs and VAEs are the most used generative models [66, 67, 52]. The combination of GAN and VAE to leverage both stochastic and realistic generation is explored by [26, 18]. To utilize the direct optimization of data likelihood of normalizing flow architectures, [25] proposed VideoFlow model that outperforms state of the art VAE models in diverse possible video generation from a sequence of past observations. A trajectory forecasting model, diversity sampling for flow (DSF) that learns a joint distribution over multiple samples in the latent space of an NF has been developed in [31] for improving both the quality and diversity of samples covering all possible modes of trajectories. While these latent variable-based methods enhance diversity, they often struggle to maintain high visual fidelity across generated frames [37]. In this work, we have used latent space sampling from VAE and NF based generative time series models, namely VRNN and GRU-NF for generating diverse video samples. We use metrics that jointly quantify diversity and reconstruction error wrt the ground truth in order to evaluate the ability of different generative models to maintain visual fidelity while also producing diverse predictions.
@@
Video Motion Transfer: Given the growing demand for high-quality interactive and immersive video experiences, particularly among users with limited bandwidth or connectivity, keypoint-based motion transfer models have emerged as an efficient solution for real-time object reconstruction and animation. These models enable a significant reduction in bandwidth requirements while maintaining video quality. Facial expressions and movements can be transferred between individuals using landmarks or keypoints combined with initial face embeddings. The Neural Talking Heads [71] algorithm leverages meta-learning on extensive video datasets to generate talking head models from just a few or even a single image of a person using adversarial training. However, this approach struggles with landmark adaptation, leading to degraded performance when landmarks are extracted from different individuals. These shortcomings are addressed by introducing a keypoint-based model that synthesizes talking head videos by integrating a source image for appearance with a driving video for motion in [62]. This method allows seamless video conferencing while minimizing bandwidth consumption without noticeable degradation in video quality. Monkey-Net proposed in [50] pioneered object-agnostic deep learning for motion transfer by enabling the animation of a source image using keypoints which are extracted from driving videos in a self-supervised manner. However, Monkey-Net relies on a zeroth-order keypoint model, leading to lower video quality when significant pose changes occur. To overcome this limitation, the First Order Motion Model (FOMM) [51] was introduced, which decouples appearance and motion information using local affine transformations and self-learned keypoints. Expanding on these models, an end-to-end unsupervised framework developed in [72] is tailored to handle substantial pose differences between source and driving images. This approach incorporates thin-plate spline motion estimation to enhance optical flow adaptability and utilizes multi-resolution occlusion masks to reconstruct missing features more effectively, resulting in higher-quality image synthesis.

### v2

@@ Abstract
Motion Transfer is an Artificial Intelligence (AI) technique that synthesizes videos by transferring motion dynamics from a driving video to a source image. In this work we propose a deep learning-based framework to enable real-time video motion transfer which is critical for enabling bandwidth-efficient applications such as video conferencing, remote health monitoring, virtual reality interaction, and vision-based anomaly detection. This is done using keypoints which serve as semantically meaningful, compact representations of motion across time, and are extracted from every video frame via a self-supervised detector. To enable bandwidth savings during video transmission we perform forecasting of keypoints using two generative time series models—Variational Recurrent Neural Networks (VRNN) and Gated Recurrent Units with Normalizing Flows (GRU-NF)—enabling both single and diverse future prediction modes. The predicted keypoints are transformed into realistic video frames using an optical flow-based module paired with a generator network, thereby facilitating accurate video forecasting and enabling efficient, low-frame-rate video transmission. Based on the application this allows the framework to either generate a deterministic future sequence or sample a diverse set of plausible futures. Experimental results across three benchmark video datasets using state-of-the-art quality and diversity metrics for video animation and reconstruction tasks demonstrate that VRNN achieves the best point-forecast fidelity (lowest MAE) in the majority of evaluated settings in applications requiring stable and accurate multi-step forecasting (e.g., video conferencing, remote patient monitoring) and is particularly competitive in higher-uncertainty, multi-modal settings. This is achieved by utilizing the superior reconstruction property of the Variational Autoencoder and by introducing recurrently conditioned stochastic latent variables that carry past contexts to capture uncertainty and temporal variation. On the other hand the GRU-NF model enables richer diversity of generated videos while maintaining high visual quality to better support tasks like AI-driven anomaly detection. This is realized by learning an invertible, exact-likelihood mapping between the keypoints and their latent representations which supports rich and controllable sampling of diverse yet coherent keypoint sequences. Our work lays the foundation for next-generation AI systems that require real-time, bandwidth-efficient, and semantically controllable video generation, with broad implications for communication, health, and manufacturing applications.
Keywords Keypoint-based motion transfer $\cdot$ Bandwidth reduction $\cdot$ Anomaly detection $\cdot$ Variational Recurrent Neural Network (VRNN) $\cdot$ Gated Recurrent Unit-Normalizing Flow (GRU-NF) $\cdot$ Diversity vs. fidelity
@@ 1 Introduction
With the growing demand for ubiquitous services, telecommuting, immersive visual interactions, and realistic simulations, real-time video based applications such as video conferencing, augmented reality (AR), virtual reality (VR) gaming, and remote medical monitoring are becoming increasingly widespread [Chan2019, Wang2021, Yang2022]. These applications demand not only high-quality visual outputs and seamless user interaction but also low-latency processing and efficient bandwidth usage—challenges that conventional video streaming and compression techniques struggle to meet effectively. Video motion transfer [Wang2021], an emerging Artificial Intelligence (AI) technique presents an effective and promising approach for realizing such applications by transferring motion from a driving video to a source frame or video, thereby enhancing user experiences and enabling more creative and engaging interactions. A key challenge in this space lies in modeling and predicting fine-grained motion or object dynamics in a way that preserves both temporal coherence and visual fidelity, while also being bandwidth-efficient for deployment in real-time, edge computing environments. Pixel-based methods if used for prediction can suffer from high computational cost, error accumulation, and limited generalization in dynamic and diverse environments. Recent advances in representation learning and generative modeling have started to address these limitations by exploring latent spaces, structured motion representations, and probabilistic forecasting [Oprea2020]. Among various representation learning based techniques for video synthesis and prediction, keypoint-based methods stand out by extracting keypoints from source and driving videos for enabling precise motion transfer [siarohin2019animating, siarohin2019first] while achieving superior video quality and compression beyond traditional codec-based methods. This not only reduces computational and bandwidth overhead, but also enables fine-grained control over motion dynamics, making keypoint-based methods ideal for real-time applications as compared to competing approaches which also claim to perform motion synthesis with high fidelity [Pondaven2024].
A critical advantage of using keypoint-based motion transfer architectures is their ability to provide a natural, low-dimensional and interpretable structure for incorporating temporal prediction using generative time series models [Chung2015, Rasul2020], thereby enabling generalization across a variety of subjects and contexts. Unlike dense optical flow fields [lu2023transflow], which require significant storage and processing per frame, keypoints offer a sparse yet semantically meaningful abstraction of motion that reduces bandwidth and computational complexity. These are key system requirements for our targeted real-time applications (video conferencing, telehealth, Virtual Reality (VR) gaming, and manufacturing industry) where inference must run on resource-constrained edge platforms. Furthermore, while explicit pose parameters [han2024prototypical] focus on specific human joints or body parts, keypoints detected in our pipeline are object-agnostic and self-supervised, allowing for broader applicability across human and non-human datasets. In contrast to pixel-level approaches, which suffer from error accumulation and sensitivity to occlusion or background clutter, keypoints focus on salient motion-driving features, enabling more stable and interpretable temporal modeling. By combining keypoint-based representation learning with deep generative forecasting models, real-time motion transfer frameworks can serve as the foundation for bandwidth-aware video AI systems, some examples of which are provided below.
@@
  • Video Conferencing: In streaming applications such as video conferencing, keypoint-based motion transfer enables efficient future frame prediction by transmitting keypoints of initial frames from a source mobile or client, allowing the pipeline to predict and generate future frames at the receiving end [Wang2021]. Such an architecture not only reduces the dependency on ubiquitous network connectivity, but also optimizes computational efficiency and bandwidth utilization.
@@
  • Telehealth: For certain critical healthcare conditions it is essential for healthcare providers along with supporting personnel such as interns to monitor patients remotely. However, in such cases patients may be hesitant to reveal their identity in virtual settings [Yang2022] and this could also be mandated by regulatory restrictions. Keypoint-based motion transfer enables future frame forecasting with anonymization that ensures both data privacy and accurate transmission of critical body movements of the patient for effective remote monitoring.
@@
  • Manufacturing Industry: A key technical challenge when deploying Artificial Intelligence (AI)-driven inspection systems in manufacturing is the scarcity of comprehensive real-world datasets that include diverse defect types. Such limitations hinder the effectiveness of machine learning models, potentially resulting in missed detections or false negatives if predictions do not correspond precisely to known defects. In vision-based anomaly detection, by forecasting future video frames of units on the production line, the system can anticipate and visualize the onset of potential defects in upcoming frames. Generating multiple plausible future sequences instead of a single prediction increases the chances of capturing visual patterns that may correspond to defects. This approach facilitates the generation of early and reliable defect warnings during real-time inspections. Moreover, the synthetic defect datasets generated through this approach expand the training data available for downstream classification or defect detection algorithms, enhancing their exposure to a wider variety of defect scenarios and improving overall robustness and detection accuracy.
@@
In this work we go beyond pre-existing keypoint-based motion transfer frameworks such as the First Order Motion Model (FOMM) [siarohin2019first] which provide spatial compression and focus on integrating probabilistic sequential generative models for keypoint forecasting in the motion transfer pipeline for realizing both spatial and temporal compression. In previous work [Bai2024], this has been demonstrated for applications involving video prediction using a Variational Recurrent Neural Network (VRNN) [Chung2015] integrated within the FOMM pipeline. In our current work we substantially expand on these preliminary findings and also investigate other generative models including Normalizing Flow and Gated Recurrent Unit-Normalizing Flow (GRU-NF) [Rasul2020] for keypoint forecasting within the motion transfer pipeline for both video prediction as well as diverse sample generation tasks. In the latter case we perform a systematic study of video diversity versus quality using a variety of metrics for measuring differences in generated realizations as well as their perceptual quality. Generative time series models are crucial here: unlike deterministic GRUs that typically assume simple output noise and collapse to a single trajectory, VRNNs introduce recurrently conditioned stochastic latents that carry history and uncertainty, while GRU-NF conditions an expressive flow on the recurrent state to model multi-modal futures without losing temporal coherence. By contrast, frame-wise VAEs/NFs without an autoregressive backbone ignore cross-time dependencies, whereas history-conditioned models such as VRNN and GRU-NF enable calibrated, diverse multi-step forecasts in multivariate settings.
@@
Our enhanced architecture in this case can provide upto 20x bandwidth savings which is over the 10x savings that can be realized from existing keypoint-based motion transfer pipelines for video reconstruction and animation [siarohin2019animating, siarohin2019first, Wang2021, Zhao2022]. To the best of our knowledge our work is the first systematic exploration of the potential of Normalizing Flow (NF) based architectures for predictive inference within a keypoint-based motion transfer framework. The rest of the paper is organized as follows. Section 2 discusses related works on video prediction, motion transfer and diverse sample generation. Section 3 describes our proposed real-time motion transfer architecture and the different forecasting models used in our work. The evaluation metrics used in this work to assess the performance of the models are explained in Section 4. Section 5 analyzes the results of our simulations based on the evaluation metrics for estimating the accuracy of generated videos and Section 6 provides a comparative analysis of the keypoint predictor models used. Finally, in Section 7 we present our conclusions and our plans for future work.
@@ 2 Related Works
Single Sample Video Prediction: The prediction and synthesis of videos that accurately represent dynamic object movements is crucial for numerous practical applications. For this purpose various video prediction techniques including direct pixel synthesis, narrowing the prediction space to high-level feature spaces [Oprea2020] have been explored by researchers. In order to accurately predict and synthesize videos by modeling object dynamics at the pixel level, numerous feature learning strategies have been explored, including adversarial training [Luc2020] and gradient difference loss functions [Mathieu2015]. Various recurrent and convolutional networks such as Convolutional Long-Short Term Memory (ConvLSTM) [Shi2015], E3d-LSTM [Wang2018], and MSPred [Villar-Corrales2022] are used for future frame prediction in raw pixel space. The common challenge faced by recurrent networks is vanishing gradient that hinders them to forecast in long term [Gao2022]. To address this, transformer-based models incorporating self-attention mechanisms have been explored for long-term video prediction [Shi2022, Ye2023, Lu2022]. While transformers have demonstrated strong performance on large text and image datasets, they are not yet well-optimized for fine-grained spatial-temporal modeling [Ming2024]. To mitigate these challenges, a simpler video prediction model entirely based on convolutional neural networks (CNN), which outperformed both recurrent neural networks (RNNs) and Vision Transformers (ViTs) has been proposed in [Gao2022]. The use of CNNs in video frame prediction has been explored by creating different types of models, i.e., Gradient Difference Loss (GDL) [Mathieu2015], cascade convolution (PredCNN) [Xu2018], and spatially-displaced convolution (SDC-Net) [Reda2018]. However, CNN-based frame synthesis is not suitable for spatially shifting pixels in input frames and is also inefficient for forecasting in large data sets [Ming2024]. Generative Adversarial Network (GAN) models have also been used to predict a sequence of future frames based on a sequence of initial video frames using pixel space as input [Aigner2018, Jang2018]. Additionally, flow-based architectures have been explored as an alternative that enables exact maximum likelihood learning utilizing invertible transformations. To perform forecasting, latent embeddings of video frames instead of raw pixels are used and future frames are reconstructed through an invertible mapping in [Pottorff2019]. The key finding here is that NF based models outperform GANs that are based on adversarial training.
@@
Inspite of their proliferation, pixel-based techniques face considerable challenges, especially for tasks requiring precise future predictions involving small or slow-moving objects. In addition, pixel-level prediction methods commonly accumulate significant errors due to inherent variability between consecutive frames. To address these challenges associated with high-dimensional pixel spaces, researchers have increasingly turned towards representation learning methods, such as semantic segmentation, human pose estimation, and keypoint-based representations. Before the widespread adoption of deep neural networks, traditional approaches employed Principal Component Analysis (PCA) or pose parameters for reconstructing video sequences with limited data, such as face boundaries [Lopez1995, Koufakis1999]. Leveraging the combined strengths of VAEs and GANs, Walker et al. [Walker2017] proposed an architecture that uses VAE to model possible future human poses in pixel space and then predicting future frames using a GAN with future poses used as conditional information. However, GANs often produce blurry and temporally inconsistent outputs [Ming2024]. Similarly, VAEs struggle to model complex video dynamics as a result of collapsing of posterior distribution to the prior [Zhou2021]. Overall, pose-guided prediction methods, despite showing promise, have largely been restricted to videos containing human subjects [Tang2019, Villegas2018, Villegas2017, Walker2017]. Semantic segmentation is another representation learning technique which instead of using raw pixel data extracts semantic elements by segmenting current frames and predict future optical flows through temporal models like Long Short-Term Memory (LSTM) networks [Ranzato2014, Terwilliger2019, Hochreiter1997]. Keypoint-based approaches for representation learning, on the other hand, explicitly represent object-level dynamics, and can provide superior outcomes for trajectory prediction and action recognition. Recent research has demonstrated the efficacy of keypoints combined with reference frames for reconstructing future video frames, thereby substantially reducing accumulated errors inherent in pixel-space forecasting [Minderer2019]. In the context of using latent space embeddings, some recent works tackle the prediction challenge by using advanced modeling techniques instead of relying on basic linear motion predictions derived from past frames [Ming2024]. In [Esser2021], a CNN-based vector quantization approach is introduced that discretizes images into a compact latent representation, enabling transformer-based autoregressive modeling for high-resolution image synthesis. Compressing high resolution videos into latent variables by using a vector quantized variational autoencoder (VQ-VAE) allowing for high resolution predictions has been proposed in [Walker2021]. However, both the CNN and VQ-VAE based models often require extensive computational resources and substantial training data. Taking these limitations into account, our work employs keypoint-based prediction methodologies, integrating these representations effectively into video reconstruction and animation in a real-time motion transfer pipeline across multiple diverse datasets.
@@
Diverse Sample Video Prediction: Incorporating uncertainty is another category of video prediction methods that allows generating diverse outcomes from a single input [Oprea2020]. Video frames form multi-modal distributions in a high dimensional space and given this level of complexity, a range of approaches have incorporated latent variables into deterministic models e.g., LSTM and CNN to introduce stochasticity and variation in predictions. Gaussian Processes combined with LSTM have been used to enable the generation of diverse samples by encoding prior knowledge of future states from past observations, allowing for a more structured distribution over possible outcomes [Shrivastava2021]. Latent variables have also been introduced into convolutional autoencoders to model video unpredictability through generative functions that capture diverse frame reconstructions [Goroshin2015]. Similarly, convolutional recurrent networks have been trained to represent future stochasticity and generate varied sequences by sampling from learned latent distributions [Hu2020]. Stochastic neural networks extend deterministic models by integrating Gaussian latent variables with convolutional encoders and spatial transformers, enabling the synthesis of multi-modal futures with enhanced diversity [Fragkiadaki2017]. The stochastic variational video prediction framework advances this direction by extending the convolutional dynamic neural advection (CDNA) model to include latent variables, allowing for different possible futures per input sample [Babaeizadeh2017]. Stochastic video generation model further improves uncertainty modeling by introducing a time-varying prior within an LSTM autoencoder, which better captures temporal dependencies in video sequences [Denton2018]. Another common approach for generating diverse predictions is by sampling from the latent space of generative models such as VAE, GAN, and NF that can learn rich probabilistic representations of data [Oprea2020]. To explore diverse video frame outcomes, GANs and VAEs are the most used generative models [Xue2016, Yang2019, Sonderby2016]. The combination of GAN and VAE to leverage both stochastic and realistic generation is explored by [Lee2018, Gur2020]. To utilize the direct optimization of data likelihood of normalizing flow architectures, [Kumar2019] proposed VideoFlow model that outperforms state of the art VAE models in diverse possible video generation from a sequence of past observations. A trajectory forecasting model, diversity sampling for flow (DSF) that learns a joint distribution over multiple samples in the latent space of an NF has been developed in [Ma2020] for improving both the quality and diversity of samples covering all possible modes of trajectories. While these latent variable-based methods enhance diversity, they often struggle to maintain high visual fidelity across generated frames [Oprea2020]. In this work, we have used latent space sampling from VAE and NF based generative time series models, namely VRNN and GRU-NF for generating diverse video samples. We use metrics that jointly quantify diversity and reconstruction error w.r.t. the ground truth in order to evaluate the ability of different generative models to maintain visual fidelity while also producing diverse predictions.
@@
Video Motion Transfer: Given the growing demand for high-quality interactive and immersive video experiences, particularly among users with limited bandwidth or connectivity, keypoint-based motion transfer models have emerged as an efficient solution for real-time object reconstruction and animation. These models enable a significant reduction in bandwidth requirements while maintaining video quality. Facial expressions and movements can be transferred between individuals using landmarks or keypoints combined with initial face embeddings. The Neural Talking Heads [Zakharov2019] algorithm leverages meta-learning on extensive video datasets to generate talking head models from just a few or even a single image of a person using adversarial training. However, this approach struggles with landmark adaptation, leading to degraded performance when landmarks are extracted from different individuals. These shortcomings are addressed by introducing a keypoint-based model that synthesizes talking head videos by integrating a source image for appearance with a driving video for motion in [Wang2021]. This method allows seamless video conferencing while minimizing bandwidth consumption without noticeable degradation in video quality. Monkey-Net proposed in [siarohin2019animating] pioneered object-agnostic deep learning for motion transfer by enabling the animation of a source image using keypoints which are extracted from driving videos in a self-supervised manner. However, Monkey-Net relies on a zeroth-order keypoint model, leading to lower video quality when significant pose changes occur. To overcome this limitation, the First Order Motion Model (FOMM) [siarohin2019first] was introduced, which decouples appearance and motion information using local affine transformations and self-learned keypoints. Expanding on these models, an end-to-end unsupervised framework, Thin Plate Spline model (TPS) developed in [Zhao2022] is tailored to handle substantial pose differences between source and driving images. This approach incorporates thin-plate spline motion estimation to enhance optical flow adaptability and utilizes multi-resolution occlusion masks to reconstruct missing features more effectively, resulting in higher-quality image synthesis. The main limitations of FOMM arise under severe occlusions, lighting shifts, and highly non-rigid motion, where locally affine warps around sparse keypoints can underfit geometry. Although the TPS [Zhao2022] addresses this, the added enhancements make the model heavier than FOMM’s locally affine pipeline, which can reduce headroom for real-time budgets on edge processors. LivePortrait [guo2024liveportrait] enhances keypoint based animation with landmark-guided keypoints and large-scale training; however, it is portrait-specific (the pipeline works with faces only) and is not purely unsupervised like FOMM since it relies on landmarks. Continuous Piecewise Affine based (CPAB) [wang2024continuous] uses a diffeomorphic continuous piecewise-affine motion space and introduce foundation model [bommasani2021opportunities] based training enhancements such SAM [kirillov2023segment] guided keypoint semantics and DINO [tumanyan2022splicing] based structure alignment, which improves robustness to complex, out-of-distribution motion. However during inference the model requires the solution of global CPAB parameters from keypoints, making motion estimation computationally costlier than FOMM and therefore harder to operate on real-time platforms.